OpenStudy (anonymous):

1. Consider the numbers 3/10 and 4/13 . a. Compare the numbers. b. Find a rational number between the two given numbers.

OpenStudy (anonymous):

You can compare by making denominators same of both the fractions..

OpenStudy (anonymous):

\[\frac {3}{10} and \frac {4}{13}\] \[\huge \rightarrow \frac { 3 \times 13}{10 \times 13} , \frac { 4 \times 10}{13 \times 10}\] \[\huge \rightarrow \frac {39}{130} , \frac {40}{130}\] \[\huge \rightarrow \frac {39}{130} < \frac {40}{130} \rightarrow \frac {3}{10} < \frac {4}{13}\]

OpenStudy (anonymous):

@jessicalove45

OpenStudy (phi):

to find a number in between you could multiply 39/130 by 2/2 and 40/130 by 2/2 to get 78/260 and 80/260 now it is easy to pick a fraction between those two.